Could you please show step by step how to multiply this out?
Thanks so much!
(x+5)(75/x - .50) = 75

Answer by stanbon(75887)

(x+5)(75/x - .50) = 75
(x)(75/x) - (x)*0.50 + 5(75/x) - 5(0.50) = 75
----
75 - 0.5x + 375/x - 0.25 = 75
---
-0.5x + 375/x - 0.25 = 0
---
Multiply thru by -x to get:
0.5x^2 - 375 + 0.25x = 0
----
Multiply thru by 100 and rearrange to get:
50x^2 + 25x - 37500 = 0
---
Divide thru by 25 to get:
2x^2 + x - 1500 = 0
----
Solve using the Quadratic formula.
----
Cheers,
Stan H.